### v5

@@ 2.3 NO’s two pillars: CO (32 transition matrix) and FO (FT pairs OTFs)
which is a transcendental equation of with being provided by various even-spectrum( biquadratic) LCO models[74, 75, 76, 77, 78, 79, 80, 81, 1, 5, 20]. Eq. 3 can be solved through direct iteration (see Method), Newton’s iteration[241], or other methods (see Discussion and Supplementary Note 6.2), with convergence typically occurring within two iterations when anisotropy is weak and the N.A. is small.
The eigenvalues and eigenvectors (i.e., eigen-polarization states) of non-uniform LFCO can be obtained by respectively substituting obtained through Eq. 3 into
and in Method.
Then, as depicted in Fig. 6 and Method, the eigenvalue pairs construct the propagation matrix , while the eigenvector pairs form the eigen-polarization state matrices and . A sandwich multiplication of these three matrices yields the 32 transition matrix (see Supplementary Note 5)
which structurally resembles the matrix exponential after Jordan decomposition (see Supplementary Note 1), realizing the transition from the two-component vector field on the input plane at to the three-component vector field on the output plane at through . Coupled with Fourier transform(FT) pairs (defined in Eqs. (S4, S5)) , Eq. 5 allows us to study the evolution and 3D distribution of the vector electric field in arbitary anisotropic dielectrics in real space, as demonstrated in Fig. 6.
Apart from CO’s in-crystal 32 transition matrix in Eq. 5 and FO’s FT pairs in Fig. 6, this LFCO model and the full trilogy, make broad use of self-built out-of-crystal optical transfer functions(OTFs) for lenses (Figs. 1, 2, 3, 4, 7 and 8 and Fig. S9), objectives (Figs. 8 and 3), half- and quarter-wave plates (Fig. 3), q-plates (Fig. 3), linear polarizers/analyzers (Fig. S11), etc., all within a custom FO framework.
@@
The field itself varies over space (2D reciprocal space, for eigensystems in Fig. S8; 3D real space, for wave diffraction in Fig. 7). This work first explores the field’s adiabatic evolution(AE) (Fig. S8 and Fig. 7) or dynamic evolution (Fig. S9 and Fig. 8) under parameter shifts, using the proposed M-M TC( M- TC in this work), implying a “change of the change”.
Arbitrary complex dielectric tensors are typically associated with birefringent-chiral-dichroic dielectrics[153, 279, 280]. In order to relate the mathematical properties of the matrix to the optical/physical properties of the material, we propose the M-M TC in Fig. 7 and Fig. S8.
@@
The potentially complex in Eq. 2 is generally non-diagonalizable in 7 out of 11 cases, unless is a normal matrix[107], corresponding to the 4 possible combinations, that is, OA + Bi, LD + CD, Bi + LD, OA + CD, i.e. two endpoints of the M-M TC’s 4 edges ‘H’, ‘!H’, ‘S’, ‘!S’, respectively.
Leveraging the reciprocal-space M-M TC in Fig. S8, we analyze the pairwise competitions among CD, OA, and LD, and how it affects of the eigensystem pairs’ distribution in 2D domain in Supplementary Note 7 (too technical for most readers from NO). We established a one-to-one extension of the three principal concepts proposed by Berry & Dennis[5] — optical singularities(i.e. EPs or singular axes), L lines, and C points — by (1) generalizing the finite set of eight optical singularities into infinite families with disk-like, annular, and crescent geometries; (2) extending L lines into L lakes and their contours, namely the L shorelines; and (3) broadening the haunting theorem associated with C points from a narrow interpretation (restricted to varing LD) to a general one (also valid for varing CD). We formulated criteria for the equivalences CD OA, OA LD, and CD LD. Ultimately, all of these confirm that eigensystems of any material are computable under our LFCO framework, foreshadowing general field reconstructions in 3D domain.
Having mastered the computation of eigensystems within arbitrary materials, the next task is to batch-reproduce camera-recorded data in real space. For this purpose, both the in-crystal CO eigenmodes with their 32 transition matrix (Eqs. 5 and 6) in reciprocal space and specific out-of-crystal FO OTFs are essential.
@@
We now revisit the real-space M-M TC in Fig. 7, upon completing the quantitative validation in Fig. S9 around the optic axis, i.e. the diabolic point of a typical biaxial material, where the eigenvalue degeneracy does not extend to the eigenvectors.
@@
Perform a counterclockwise scan along the three edges of the M-M TC. First, increasing OA from 0 along the ‘H’ edge of the M-M TC yields the AE diagram of the optical rotation field from vertice ‘Bi’ to ‘OA’ in Fig. 7e (pointed by the light blue arrows), ultimately converging to Bloembergen et al.’s experiment: chiral CR[159] through -HIO3. Next, while maintaining OA, as LD increases along the ‘Im’ edge from 0 (denoted by the light green arrows), Brenier’s experimental results on chirality versus dichroism[160] of acentric chiral Nd3+-doped BZBO are obtained, with a clear criterion for OA LD. That is, when pump’s polarization is aligned parallel to the polarization of Voigt wave while the analyzer is set perpendicular, central extinction is observed in the far field. Following this, by canceling OA and keeping only LD, the angular absorption distribution of laser crystal KGd(WO4)2 is obtained[139, 81]. Then, by gradually reducing LD along the ‘S’ edge (antiparallel to the light red arrows), the Pancharatnam phenomenon[158] emerges.
Finally, as LD is reduced to 0, only Bi remains, including uniaxiality (Fig. S10b and Fig. S11), biaxiality (Fig. S10a and Fig. S11b1), and their hyperbolic counterparts (Fig. S10a,b). As an effect caused by the off-diagonal elements of the 22 eigen-polarization matrix and an embodiment of conservation law, the global grasp and detailed calculation of SOI play a significant role in not only LCO[215, 216, 217, 219, 221, 223, 222, 224, 225, 226] (see Fig. 6 and Figs. S9, S11 and SA3-SA5) but also NCO[247, 281, 6, 255] (see Fig. 3).
@@ 2.5 Tightly focused light in highly anisotropic materials
In the extreme case of strong linear interaction between highly anisotropic materials and tightly focused light fields — an area at the forefront of both industry and academia in laser processing, aberration correction, and inverse focal engineering — our LFCO model provides a unified solution, as shown in Fig. 8. Here, we show the underlying mechanism of Raman spike[161] (Fig. 8b) in CR, the reverse engineering technique for the targeted vector complex field at the focal plane free of Zernike polynomial (Fig. 8c1), the “Cherenkov cone” outside the inner beam (Fig. 8c2) marking the computational boundary, the R-L and o-e decomposition analysis for high-N.A. pump (Fig. 8c4), and double conical refraction and the optical field self-twisting effects it induces (Fig. 8d1).
@@
Using the crytal-version vector angular spectrum method(ASM), i.e. this LFCO model, without any optimization, we explored various tightly focused field distributions while operating at the edge of the sampling theorem, with each subfigure (except Fig. 8c2) in Fig. 8 nearing the computational limits just before aliasing errors occur. Initially, we believed that the origin of the outer cone in Fig. 8c2 was physical because: (1) it began forming even before propagation, and (2) it appeared at the center of the image rather than being reflected from the edges. However, we later discovered that it was still caused by aliasing errors originating from circular convolution. This conclusion was drawn because, in order to shift the focal plane of Fig. 8c2 inside the LN materiral, we set the incident field distribution at the material’s entrance to a LP Gaussian beam with a N.A. close to 1 (see the top row of Fig. 8c3), propagating backward by mm. However, when was set to zero, the outer cone disappeared. In contrast, Fig. 8c4 uses the transfer function of the objective lens instead of backward propagation to shift the focal plane below the upper surface of LN, completely eliminating the aliasing error.
We find that by setting negative, thereby transforming the refractive index surface from an ellipsoid to a dielectric-type two-sheet hyperboloid[114, 118, 119, 113, 122], the two optical axes can be brought closer together, as shown in Fig. 8d2. For KTP crystal at 1064 nm, the angle between the two optical axes decreases from 34.6 in the ellipsoidal case to 7.7 in the hyperboloidal case, allowing us to compute the full evolution inside and outside the crystal in a non-paraxial manner over relatively longer propagation distances of the entire crystal-2f system. Optical field knotting in the 1st and 3rd column of Fig. 8d1 are beyond our present mathematical understanding and techniques for explanation.
@@ 2.6 Superstructure of this LCO model: NCO QNCO
Both the NFs and the pumps involved in the interaction, are constrained by their own monochromatic passive LCO wave equation, necessitating independent diffraction as the crystal’s eigenmodes (Fig. 5d,e,g,h).
The NCO parametric frequency conversion process must first satisfy energy conservation, followed by momentum conservation (often described as wave vector or phase matching), which further tests the precision of LCO eigenvalue calculations (Fig. 5d,e). If the anisotropy of the second-order nonlinear coefficient tensor is additionally involved, the accurate computation of eigenvectors (Fig. 5g,h) in the frame[277, 13, 3, 4] within LCO must also be ensured as a prerequisite.
@@
Having conducted large-scale numerical validation of known NCO phenomena in Section 2.2, thereby indirectly substantiating the LCO model advanced in this paper, it is imperative to further establish the model’s forward-looking predictive capabilities, for the potentially vast range of yet-unknown NCO phenomena.
To this purpose, we propose simulations of spontaneous parametric down-conversion(SPDC) in 1D, 2D, and 3D NPCs in Fig. 9 (and Fig. SB6), as a preliminary step toward developing a comprehensive theoretical and mathematical FO framework for quantum NCO(QNCO) nonlinear quantum optics(NQO) QO.
@@ 3 Discussion
Conversely, numerically stable models that allow for the treatment of thick non-Hermitian slabs fail entirely to resolve EPs. These models share the same structure of classical plane wave , whose two distinct forms based on real spherical/rectangular coordinates, with different independent variables , both satisfy the same wave/characteristic Eq. 2. All models choosing as input variable typically requires numerical solutions for quartics, while the rest models depends on possess simple closed-form solutions to biquadratics. However, the former satisfies FT and boundary conditions, while the latter does not.
@@
By integrating the FO transfer functions of our custom-developed optical instruments, our trilogy (this LFCO model subsequent NFCO models) has also successfully reproduced and further explored numerous complicated experimental results in both LCO and NCO, unveiling the unified mathematical and physical essence underlying these diverse phenomena. This forward LFCO model is also inversely applied for inverse focal design in high N.A. laser writing. Lastly, we offer some initial demonstrations of the potential future applications of this powerful LCO model in (quantum) NCO, including spontaneous parametric down-conversion(SPDC) in 3D nonlinear photonic crystals(NPCs).
Such a novel paradigm (the M-M TC and FO OTFs with crystal-2f configuration) and the newly discovered - and -space phenomena will inject theoretical, experimental and computational vitality into LCO & NCO, as well as CO, FO, LO, NO, and QO, breathing new life into their future developments. It also lays a solid foundation for all upper-level architectures based on light-matter interactions and will inspire a fresh wave of rethinking in all wave theories, including quantum mechanics, elastodynamics[288, 227] and (nonlinear) (ultra)sound propagation[270, 289, 290, 291, 292, 271, 293, 294].
@@ 4.1 Boundary conditions for laboratory settings
Then the wave equation (see Supplementary Note 2), together with Eq. 6, contains all the necessary information for solving the distribution of electromagnetic field inside and outside the slab made of typical optical materials, provided that a two-dimensional(2D) distribution of the vector pump (’s transverse components ) in front of the slab () is given, which means neither divergence equations[191, 299, 231, 240, 91, 90, 300, 301] nor boundary conditions[191, 299] are even required.
@@ 4.2 Eigensystem corrections for phase continuity
where both the input real ray direction for uniform eigenvalue and the output non-uniform eigenvalue are ultimately functions of , and mutually coupled, as illustrated in Fig. 10a2.
One can unravel this coupling by first discarding the imaginary part of the complex non-uniform wave vector in Eq. 8b, retaining only the real part ’s unit vector from Eq. 9a as the ray direction in space. The input variable is then inserted into Eq. 9b to form a transcendental equation which asymptotically converges to the ground true eigenvalue of the booker quartic on condition that
where material absorption is sufficiently low with typically when a transmission spectrum is still present, as exemplified by Brenier’s laser crystal[139, 81] in Fig. 7b,e.
Berry & Dennis’s input variable can be obtained from Fig. 10a2 as an intermediate data. Inserting it into Fig. 10b yields Berry & Dennis’s eigensystem , where the eigenvector also converges to the ground truth under the condition of Eq. 10 (see Supplementary Note 2.2), after undergoing the operation shown in Fig. 10c.

### v6

@@ 2.3 NO’s two pillars: CO (32 transition matrix) and FO (FT pairs OTFs)
which is a transcendental equation of with being provided by various even-spectrum( biquadratic) LCO models[74, 75, 76, 77, 78, 79, 80, 81, 1, 5, 20]. Eq.˜3 can be solved through direct iteration (see Method), Newton’s iteration[241], or other methods (see Discussion and Supplementary Note 6.2), with convergence typically occurring within two iterations when anisotropy is weak and the N.A. is small.
The eigenvalues and eigenvectors (i.e., eigen-polarization states) of non-uniform LFCO can be obtained by respectively substituting obtained through Eq.˜3 into
and in Method. Then, as depicted in Fig.˜6 and Method, the eigenvalue pairs construct the propagation matrix , while the eigenvector pairs form the eigen-polarization state matrices and . A sandwich multiplication of these three matrices yields the 32 transition matrix (see Supplementary Note 5)
which structurally resembles the matrix exponential after Jordan decomposition (see Supplementary Note 1), realizing the transition from the two-component vector field on the input plane at to the three-component vector field on the output plane at through . Coupled with Fourier transform(FT) pairs (defined in Eqs. (S4, S5)) , Eq.˜5 allows us to study the evolution and 3D distribution of the vector electric field in arbitary anisotropic dielectrics in real space, as demonstrated in Fig.˜6.
Apart from CO’s in-crystal 32 transition matrix in Eq.˜5 and FO’s FT pairs in Fig.˜6, this LFCO model and the full trilogy, make broad use of self-built out-of-crystal optical transfer functions(OTFs) for lenses (Figs.˜1, 2, 4, 3, 7 and 8 and Fig. S9), objectives (Figs.˜8 and 4), half- and quarter-wave plates (Fig.˜4), q-plates (Fig.˜4), linear polarizers/analyzers (Fig. S11), etc., all within a custom FO framework.
@@
The field itself varies over space (2D reciprocal space, for eigensystems in Fig. S8; 3D real space, for wave diffraction in Fig.˜7). This work first explores the field’s adiabatic evolution(AE) (Fig. S8 and Fig.˜7) or dynamic evolution (Fig. S9 and Fig.˜8) under parameter shifts, using the proposed M-M TC( M- TC in this work), implying a “change of the change”.
Arbitrary complex dielectric tensors are typically associated with birefringent-chiral-dichroic dielectrics[153, 279, 280]. In order to relate the mathematical properties of the matrix to the optical/physical properties of the material, we propose the M-M TC in Fig.˜7 and Fig. S8.
@@
The potentially complex in Eq.˜2 is generally non-diagonalizable in 7 out of 11 cases, unless is a normal matrix[107], corresponding to the 4 possible combinations, that is, OA + Bi, LD + CD, Bi + LD, OA + CD, i.e. two endpoints of the M-M TC’s 4 edges ‘H’, ‘!H’, ‘S’, ‘!S’, respectively.
Leveraging the reciprocal-space M-M TC in Fig. S8, we analyze the pairwise competitions among CD, OA, and LD, and how it affects of the eigensystem pairs’ distribution in 2D domain in Supplementary Note 7. We established a one-to-one extension of the three principal concepts proposed by Berry & Dennis[5] — optical singularities(i.e. EPs or singular axes), L lines, and C points — by (1) generalizing the finite set of eight optical singularities into infinite families with disk-like, annular, and crescent geometries; (2) extending L lines into L lakes and their contours, namely the L shorelines; and (3) broadening the haunting theorem associated with C points from a narrow interpretation (restricted to varing LD) to a general one (also valid for varing CD). We formulated criteria for the equivalences CD OA, OA LD, and CD LD. Ultimately, all of these confirm that eigensystems of any material are computable under our LFCO framework, foreshadowing general field reconstructions in 3D domain.
Having mastered the computation of eigensystems within arbitrary materials, the next task is to batch-reproduce camera-recorded data in real space. For this purpose, both the in-crystal CO eigenmodes with their 32 transition matrix (Eqs.˜5 and 6) in reciprocal space and specific out-of-crystal FO OTFs are essential.
@@
We now revisit the real-space M-M TC in Fig.˜7, upon completing the quantitative validation in Fig. S9 around the optic axis, i.e. the diabolic point of a typical biaxial material, where the eigenvalue degeneracy does not extend to the eigenvectors.
@@
Perform a counterclockwise scan along the three edges of the M-M TC. First, increasing OA from 0 along the ‘H’ edge of the M-M TC yields the AE diagram of the optical rotation field from vertice ‘Bi’ to ‘OA’ in Fig.˜7e (pointed by the light blue arrows), ultimately converging to Bloembergen et al.’s experiment: chiral CR[159] through -HIO3. Next, while maintaining OA, as LD increases along the ‘Im’ edge from 0 (denoted by the light green arrows), Brenier’s experimental results on chirality versus dichroism[160] of acentric chiral Nd3+-doped BZBO are obtained, with a clear criterion for OA LD. That is, when pump’s polarization is aligned parallel to the polarization of Voigt wave while the analyzer is set perpendicular, central extinction is observed in the far field. Following this, by canceling OA and keeping only LD, the angular absorption distribution of laser crystal KGd(WO4)2 is obtained[139, 81]. Then, by gradually reducing LD along the ‘S’ edge (antiparallel to the light red arrows), the Pancharatnam phenomenon[158] emerges.
Finally, as LD is reduced to 0, only Bi remains, including uniaxiality (Fig. S10b and Fig. S11), biaxiality (Fig. S10a and Fig. S11b1), and their hyperbolic counterparts (Fig. S10a,b). As an effect caused by the off-diagonal elements of the 22 eigen-polarization matrix and an embodiment of conservation law, the global grasp and detailed calculation of SOI play a significant role in not only LCO[215, 216, 217, 219, 221, 223, 222, 224, 225, 226] (see Fig.˜6 and Figs. S9, S11 and SA3-SA5) but also NCO[247, 281, 6, 255] (see Fig.˜4).
@@ 2.5 Tightly focused light in highly anisotropic materials
In the extreme case of strong linear interaction between highly anisotropic materials and tightly focused light fields — an area at the forefront of both industry and academia in laser processing, aberration correction, and inverse focal engineering — our LFCO model provides a unified solution, as shown in Fig.˜8. Here, we show the underlying mechanism of Raman spike[161] (Fig.˜8b) in CR, the reverse engineering technique for the targeted vector complex field at the focal plane free of Zernike polynomial (Fig.˜8c1), the “Cherenkov cone” outside the inner beam (Fig.˜8c2) marking the computational boundary, the R-L and o-e decomposition analysis for high-N.A. pump (Fig.˜8c4), and double conical refraction and the optical field self-twisting effects it induces (Fig.˜8d1).
@@
Using the crytal-version vector angular spectrum method(ASM), i.e. this LFCO model, without any optimization, we explored various tightly focused field distributions while operating at the edge of the sampling theorem, with each subfigure (except Fig.˜8c2) in Fig.˜8 nearing the computational limits just before aliasing errors occur. Initially, we believed that the origin of the outer cone in Fig.˜8c2 was physical because: (1) it began forming even before propagation, and (2) it appeared at the center of the image rather than being reflected from the edges. However, we later discovered that it was still caused by aliasing errors originating from circular convolution. This conclusion was drawn because, in order to shift the focal plane of Fig.˜8c2 inside the LN materiral, we set the incident field distribution at the material’s entrance to a LP Gaussian beam with a N.A. close to 1 (see the top row of Fig.˜8c3), propagating backward by mm. However, when was set to zero, the outer cone disappeared. In contrast, Fig.˜8c4 uses the transfer function of the objective lens instead of backward propagation to shift the focal plane below the upper surface of LN, completely eliminating the aliasing error.
We find that by setting negative, thereby transforming the refractive index surface from an ellipsoid to a dielectric-type two-sheet hyperboloid[114, 118, 119, 113, 122], the two optical axes can be brought closer together, as shown in Fig.˜8d2. For KTP crystal at 1064 nm, the angle between the two optical axes decreases from 34.6 in the ellipsoidal case to 7.7 in the hyperboloidal case, allowing us to compute the full evolution inside and outside the crystal in a non-paraxial manner over relatively longer propagation distances of the entire crystal-2f system. Optical field knotting in the 1st and 3rd column of Fig.˜8d1 are beyond our present mathematical understanding and techniques for explanation.
@@ 2.6 Superstructure of this LCO model: NCO QNCO
Both the NFs and the pumps involved in the interaction, are constrained by their own monochromatic passive LCO wave equation, necessitating independent diffraction as the crystal’s eigenmodes (Fig.˜5d,e,g,h).
The NCO parametric frequency conversion process must first satisfy energy conservation, followed by momentum conservation (often described as wave vector or phase matching), which further tests the precision of LCO eigenvalue calculations (Fig.˜5d,e). If the anisotropy of the second-order nonlinear coefficient tensor is additionally involved, the accurate computation of eigenvectors (Fig.˜5g,h) in the frame[277, 13, 3, 4] within LCO must also be ensured as a prerequisite.
@@
Having conducted large-scale numerical validation of known NCO phenomena in Section˜2.2, thereby indirectly substantiating the LCO model advanced in this paper, it is imperative to further establish the model’s forward-looking predictive capabilities, for the potentially vast range of yet-unknown NCO phenomena.
To this purpose, we propose simulations of spontaneous parametric down-conversion(SPDC) in 1D, 2D, and 3D NPCs in Fig.˜9 (and Fig. SB6), as a preliminary step toward developing a comprehensive theoretical and mathematical FO framework for quantum NCO(QNCO) nonlinear quantum optics(NQO) QO.
@@ 3 Discussion
Conversely, numerically stable models that allow for the treatment of thick non-Hermitian slabs fail entirely to resolve EPs. These models share the same structure of classical plane wave , whose two distinct forms based on real spherical/rectangular coordinates, with different independent variables , both satisfy the same wave/characteristic Eq.˜2. All models choosing as input variable typically requires numerical solutions for quartics, while the rest models depends on possess simple closed-form solutions to biquadratics. However, the former satisfies FT and boundary conditions, while the latter does not.
@@
By integrating the custom-developed FO transfer functions of optical instruments, our trilogy (this LFCO model subsequent NFCO models) has also successfully reproduced and further explored numerous complicated experimental results in both LCO and NCO, unveiling the unified mathematical and physical essence underlying these diverse phenomena. This forward LFCO model is also inversely applied for inverse focal design in high N.A. laser writing. Lastly, we offer some initial demonstrations of the potential future applications of this powerful LCO model in (quantum) NCO, including spontaneous parametric down-conversion(SPDC) in 3D nonlinear photonic crystals(NPCs).
Such a novel paradigm (the M-M TC and FO OTFs with crystal-2f configuration) and the newly predicted - and -space phenomena will inject theoretical, experimental and computational vitality into LCO & NCO, as well as CO, FO, LO, NO, and QO, breathing new life into their future developments. It also lays a solid foundation for all upper-level architectures based on light-matter interactions.
@@ 4.1 Boundary conditions for laboratory settings
Then the wave equation (see Supplementary Note 2), together with Eq.˜6, contains all the necessary information for solving the distribution of electromagnetic field inside and outside the slab made of typical optical materials, provided that a two-dimensional(2D) distribution of the vector pump (’s transverse components ) in front of the slab () is given, which means neither divergence equations[191, 299, 231, 240, 91, 90, 300, 301] nor boundary conditions[191, 299] are even required.
@@ 4.2 Eigensystem corrections for phase continuity
where both the input real ray direction for uniform eigenvalue and the output non-uniform eigenvalue are ultimately functions of , and mutually coupled, as illustrated in Fig.˜10a2.
One can unravel this coupling by first discarding the imaginary part of the complex non-uniform wave vector in Eq.˜8b, retaining only the real part ’s unit vector from Eq.˜9a as the ray direction in space. The input variable is then inserted into Eq.˜9b to form a transcendental equation which asymptotically converges to the ground true eigenvalue of the booker quartic on condition that
where material absorption is sufficiently low with typically when a transmission spectrum is still present, as exemplified by Brenier’s laser crystal[139, 81] in Fig.˜7b,e.
Berry & Dennis’s input variable can be obtained from Fig.˜10a2 as an intermediate data. Inserting it into Fig.˜10b yields Berry & Dennis’s eigensystem , where the eigenvector also converges to the ground truth under the condition of Eq.˜10 (see Supplementary Note 2.2), after undergoing the operation shown in Fig.˜10c.